Circular No. 6441

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996am IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6418) reports the discovery of a supernova on CCD images taken by S. Chan on the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1996am is located at α = 21^{h}46^{m}07^{s}.48, δ = −44^{o}00′12′′.7 (equinox 2000.0), which is near the nucleus of the host galaxy (itself situated near Abell 3809). A nearby star of mag 12 is at α = 21^{h}46^{m}12^{s}.57, δ = −43^{o}58′51′′.3. SN 1996am is visible on CCD images taken on July 12 at V ∼ R ∼ 19.0, and it was not detected on similar frames taken on July 3. A spectrum (range 350–1000 nm) obtained on July 25 with the MSSSO 2.3-m Imager + Grism shows that this is a type-Ia supernova near maximum. The host galaxy’s redshift, measured from Hα emission at the position of the supernova, is z = 0.065, which is coincident with the redshift of Abell 3809 (z = 0.062).

V1974 CYGNI S. Balman, J. S. Gallagher and M. Orio, University of Wisconsin at

Madison, report: “Using the Kitt Peak 3.5-m WIYN telescope (+ CCD imager) on May 27 UT, we detected an Hα nebulosity associated with the classical nova V1974 Cyg (N Cyg 1992). The faint, diffuse, Hα emission around the nova had dimensions in each direction as follows: north, 125′′; south, 190′′; west, 140′′; east, 142′′. An elliptical cavity was also prominent around the nova, extending as follows: north, 100′′; south, 21′′; west, 35′′; east, 38′′. The Hα nebulosity was found to peak in the southwest from a region 152′′ × 208′′ with long, thin, chevron-like filamentary structures. The size of the nebulosity is consistent with light travel time arguments for a reflection nebula. Our observation confirms the previous spectroscopic result of extensive Hα emission from the vicinity of the nova by Rosino et al. (1996, AAp, submitted). On June 16, the nebulosity showed no substantial changes; the narrowband Hα luminosity of the nebula in the June observation is 2–3 times that of the nova, which also showed a shell diameter of 0′′.8 with an extension to the northwest.”

COMET 57P/DU TOIT-NEUJMIN-DELPORTE This comet is in outburst; m_{1} estimates: July 24.78 UT, 13.3 (A.

Nakamura, Kuma, Japan, 0.60-m reflector + CCD); 26.08, 13.3 (H. Mikuž, Ljubljana, Slovenia, 0.20-m reflector + CCD + V filter); 26.45, 11.8: (A. Hale, Cloudcroft, NM, 0.41-m reflector, visual); 28.04, 11.9: (M. Plsek, Lelekovice, Czech Republic, 0.35-m reflector, visual).

1996 July 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green